OpenStudy (anonymous):

Sal’s Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and The profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal’s profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold.

OpenStudy (anonymous):

1. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work below:

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

2x + 3y = 1,470 3y = 1470 – 2x y = 1470/3 – 2/3x y = 490 – 2/3x y = -2/3 * x + 490 the slope is -2/3 and y intercept is 490 this is what i got

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

but i need help with this question 3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.

OpenStudy (michele_laino):

using function notation, you have to replace \(y\) with \(f(x)\) so we have: \[f\left( x \right) = - \frac{2}{3}x + 490\]

OpenStudy (anonymous):

Okay thank you. 2 more..

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

5. Suppose Sal’s total profit on lunch specials for the next month is $1,593. The profit amounts are the same: $2 for each sandwich and $3 for each wrap. In a paragraph of at least three complete sentences, explain how the graphs of the functions for the two months are similar and how they are different

OpenStudy (michele_laino):

we have to apply the same method of before, in order to write the second function

OpenStudy (michele_laino):

so we can write this: \(2x+3y=1593\) please solve for \(y\)

OpenStudy (anonymous):

uhmm y= -2x/3 + 531

OpenStudy (michele_laino):

correct! We can write: \[g\left( x \right) = - \frac{2}{3}x + 531\]

OpenStudy (michele_laino):

I called the second function with \(g(x)\)

OpenStudy (michele_laino):

so, what can you conclude?

OpenStudy (michele_laino):

here you have to compare \(f(x)\) with \(g(x)\)

OpenStudy (anonymous):

ummm im not sure :/

OpenStudy (michele_laino):

have those function the same slope?

OpenStudy (michele_laino):

functions*

OpenStudy (anonymous):

yes they do

OpenStudy (michele_laino):

correct! we have a common feature

OpenStudy (michele_laino):

have those functions different y-intercepts?

OpenStudy (anonymous):

uhmm yes they do?

OpenStudy (michele_laino):

correct! since the y-intercept of \(f(x)\) is 490, whereas the y-intercept of \(g(x)\) is 531

OpenStudy (anonymous):

yes! you are a life saver!!

OpenStudy (michele_laino):

thanks!! :) :)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (michele_laino):

:)

kward64:

is this all the same person

kward64:

bc i sure looks like it

Eiwoh2:

No, this is a question transferred over from the predecessor site, OpenStudy lol These are different users from there that got turned into “OpenStudy” users, since it dosent exist anymore.

kward64:

oh ok

Eiwoh2:

This was also 6 years ago lol

kward64:

just didnt know

kward64:

ik